An electron is moving at a speed of 1.20 ✕ 104 m/s in a circular path of radius of 2.5 cm inside a solenoid. The magnetic field of the solenoid is perpendicular to the plane of the electron's path. The solenoid has 25 turns per centimeter.

(a) Find the strength of the magnetic field inside the solenoid.
μT

(b) Find the current in the solenoid.
mA
Answer #1

a)

Given

n=25 turns/cm =25*102 turns/m

Here centripetal force is balanced by magnetic force

mv2/r = qvB

=>B=mv/qr =(9.11*10-31)(1.2*104)/(1.6*10-19)(0.025)

B=2.733*10-6 = 2.733 uT

b)

Magnetic field of a solenoid is given by

B=uonI

=>I=B/uon =2.733*10-6)/(25*102)(4pi*10-7)

I=8.7*10-4 A =0.87 mA